OpenStudy (anonymous):

solve the equation 8n-(-2)/7 =6 n=???

jimthompson5910 (jim_thompson5910):

I'm assuming that this is \[\Large 8n - \frac{-2}{7} = 6\] correct?

OpenStudy (anonymous):

sorry no the 8n-(-2) are on top together and the seven is on the bottom

jimthompson5910 (jim_thompson5910):

so it's \[\Large \frac{8n - (-2)}{7} = 6\] right?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

Ok any ideas on the first step?

OpenStudy (anonymous):

nooope haha

jimthompson5910 (jim_thompson5910):

How about we rewrite that -(-2) as +2. We can do this because subtracting a negative is just like adding So \[\Large \frac{8n - (-2)}{7} = 6\] becomes \[\Large \frac{8n +2}{7} = 6\]

jimthompson5910 (jim_thompson5910):

Next, notice how we're dividing by 7 on the left side. We want to move this 7 over to isolate n. So we have to undo this by doing the opposite What is the opposite of dividing by 7?

OpenStudy (anonymous):

multiplying?

jimthompson5910 (jim_thompson5910):

good, multiplying by what?

OpenStudy (anonymous):

7 or -7?

jimthompson5910 (jim_thompson5910):

positive 7

jimthompson5910 (jim_thompson5910):

Because the 7 in the denominator is positive

jimthompson5910 (jim_thompson5910):

So multiply both sides by 7 If you multiply the left side by 7, the 7 in the denominator will cancel. This will leave you with 8n+2 on the left side ------------------------------------------- Then multiply the right side by 7 to balance things out to get 7*6 = 42 ----------------------------------------------------------------- So \[\Large \frac{8n+2}{7} = 6\] becomes \[\Large 8n+2 = 42\]

jimthompson5910 (jim_thompson5910):

now your goal is to solve 8n+2 = 42

OpenStudy (anonymous):

5?

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

Thanks you really are helping me alot haha

jimthompson5910 (jim_thompson5910):

lol I'm very glad I am